OpenStudy (anonymous):

need help with solving systems of equations with three variables, I'm a lil confused 4x+y=-7 x-2z=4 3y+2z=8

OpenStudy (anonymous):

nvm I got it

OpenStudy (wolf1728):

A) 4x+y=-7 B) x-2z=4 C) 3y+2z=8 Add equations B & C. x + 3y = 12 Is that enough of a start for you?